4. Your roommate Pat decides to help with the party. The amounts of brownies and cake which can be produced in an hour (if all resources are devoted to that good) are shown by the following table:

                                                                               YOU                           PAT        

                        BROWNIES PER HOUR                  8                                 2

                        CAKE PER HOUR                            2                                 1

a. Who has an absolute advantage in producing brownies? Cake?

            b. For you, what is the opportunity cost of one brownie? Of one cake?

            c. For Pat, what is the opportunity cost of one brownie? Of one cake?

            d. Should you and Pat specialize? Who should produce what?

            e. Suppose that you have so many guests that you have to split up the party into two  rooms. You throw one party and Pat throws the other. What are the terms of trade?                        (At what rate would you each be willing to trade?)

f. Using graphs for both you and Pat, show that trade allows each of you to

    achieve a point on your consumption possibilities curve which is greater than  

    your production possibilities curve. Be sure to label all curves and axes.

g. What is this principle called?
Answer #1

a) Absolute Advantage:

Absolute advantage is the ability of a country to produce a good or service at a lower per unit cost as compared to any other country that produces same good or service.

YOU has an absolute advantage in the production of brownies and cake because YOU are able to produce more of both goods as compared to PAT.

b) Opportunity cost is the value of next best alternative foregone.

YOU:

8 brownies = 2 cakes

1 brownie = 2/8 = 0.25 cake

Opportunity cost of 1 brownie is 0.25 cake.

2 cakes = 8 brownies

1 cake = 8/2 = 4 brownies

Opportunity cost of 1 cake is 4 brownies.

c) PAT:

2 brownies = 1 cakes

1 brownie = 1/2 = 0.5 cake

Opportunity cost of 1 brownie is 0.5 cake.

1 cakes = 2 brownies

Opportunity cost of 1 cake is 2 brownies.

d) Comparative Advantage

A country has a comparative advantage in producing that good if the opportunity cost of producing that good is lower in that country as compared to another country.

PAT has comparative advantage in the production of cake because of lower opportunity cost in the production of cakes by PAT.

YOU has comparative advantage in the production of brownies because of lower opportunity cost.

So, YOU should produce only brownies while PAT has to produce only cakes.